# Analysis of Semi-Supervised Learning on Hypergraphs

### v1

@@ 1 Introduction
    Since the dependence on $v$ in the objective function arises solely from pairwise interactions between vertices, the resulting regularization is inherently first-order. Consequently, the hypergraph itself contributes only to defining the support of interactions—that is, which and how strongly vertex pairs are connected—not to specifying the order or nature of the regularization being applied. We show in Theorems 3.3 and 3.4 that the solution of (2), in an appropriate limit, behaves as the sum of first order energies.
@@
In this paper, we study the consistency of (2) through the lens of the latter two complementary notions of convergence. Pointwise convergence allows us to show that the operator appearing in the discrete Euler–Lagrange equation of (2) converges to a non-trivial continuum operator, which behaves qualitatively like a $p$-Laplace operator. Variational convergence, on the other hand, enables a precise characterization of the well- and ill-posedness of hypergraph learning in the semi-supervised setting, as a function of the hypergraph construction parameters. The appropriate analytical framework for convergence of minimizers is $\Gamma$-convergence [11], which requires a common metric space in which both discrete and continuum functions can be embedded. To this end, we rely on the $\mathrm{TL}^{p}$ space introduced in [37], which is based on tools from optimal transport [84, 73]. We review these methodologies in Section 2.
A notable observation is that the non-trivial continuum operator obtained via pointwise convergence differs from the operator arising in the Euler–Lagrange equation of the limiting energy obtained through variational convergence. To the best of our knowledge, this phenomenon has not been previously documented in the context of graph-based algorithms. This discrepancy highlights a subtle but important distinction between the two modes of convergence and underscores the need to consider both when analyzing the continuum limits of graph- or hypergraph-based learning models.
@@ 1.1 Contributions
    Pointwise Convergence to Continuum Operator: We establish the pointwise convergence of the discrete Euler–Lagrange operators arising in hypergraph learning to a non-trivial continuum operator that behaves like a $p$-Laplacian.
@@
    Mismatch Between Pointwise and Variational Limits: We show that the limiting operator obtained through pointwise convergence differs from the variational limit, highlighting a previously undocumented distinction between these two modes of convergence in graph-based algorithms.
@@ 1.2 Related works
On the technical side, our proofs rely on nonlocal approximation results for $\mathrm{W}^{1,p}$ initially formulated in [10] and then considered under the $\Gamma$-convergence lens in [70]. The latter were extended to the discrete-to-continuum setting in [37, 80]. For other discrete-to-continuum results approximating $\mathrm{W}^{s,2}$ we refer to [25, 90, 36]. Finally, we note that our results are purely asymptotic. Discrete-to-continuum rates have for example been obtained on similar graph learning problems in [19, 89, 27].
@@ 2.1 The $\mathrm{TL}^{p}$ Space
[37, Proposition 3.12] Let $(\mu,u)\in\mathrm{TL}^{p}$ where $\mu$ is absolutely continuous with respect to Lebesgue measure and let $\{(\mu_{n},u_{n})\}_{n=1}^{\infty}$ be a sequence in $\mathrm{TL}^{p}$. The following are equivalent:
@@
In order to compare discrete functions to continuum ones, we will let $\mu_{n}$ be the empirical measures of our samples $\{x_{i}\}_{i=1}^{n}$ and $\mu$ will be the measure from which the points are sampled. Furthermore, $\{u_{n}\}_{n=1}^{\infty}$ and $u$ will respectively be the minimizers of our discrete and continuum objectives. In order to use the above result, we need certain transport maps $T_{n}$ whose existence is guaranteed by the next theorem [41, Theorem 1.1].
@@ 3.2 Hypergraph setting
and the constants
$\sigma_{\eta}^{(k,p)}$ $=\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})|(\tilde{z}_{1})_{d}|^{p}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1},$
$\sigma_{\eta}^{(k,p,1)}$ $=\int_{(\mathbb{R}^{d})^{k}}\hskip-8.53581pt\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})|(\tilde{z}_{1})_{d}|^{p-2}(\tilde{z}_{1})_{1}^{2}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1},$
$\sigma_{\eta}^{(k,p,2)}$ $=\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})|(\tilde{z}_{1})_{d}|^{p-2}(\tilde{z}_{1})_{d}(\tilde{z}_{2})_{d}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}$
@@
where $\Delta$ denotes the regular continuum Laplacian operator.
Remark 3.1 (Relationship between $\Delta_{\infty}^{(k,p)}$ and the $p$-Laplacian operator).
For $k=1$, we have
$\Delta_{\infty}^{(1,p)}(u)(x)=\frac{\sigma_{\eta}^{(1,p)}}{2\rho(x)}\mathrm{div}(\|\nabla u(x))\|_{2}^{p-2}\nabla u(x)\rho(x)^{2})$
so that $\Delta_{\infty}^{(1,p)}$ is a weighted $p$-Laplacian operator. The key identity used in the previous computation is $\sigma_{\eta}^{(1,p,1)}(p-1)=\sigma_{\eta}^{(1,p)}$. For $k>1$, Corollary 4.2 yields an equivalence bound of the form $\frac{\sigma_{\eta}^{(1,p,1)}(p-1)}{\sigma_{\eta}^{(1,p)}}\in[c,C]$. This implies that qualitatively, up to constants, $\Delta_{\infty}^{(k,p)}$ acts like a $p$-Laplacian operator.
@@ 3.5 Variational problems for higher-order hypergraph learning
The latter set can be shown to be very closely related to the Sobolev space $\mathrm{W}^{p,2}(\Omega)$ [25, Lemma 17]. Finally, for $q\geq 1$ and positive sequences $\{\lambda_{k}\}_{k=1}^{q}\subseteq\mathbb{R}$, $P:=\{p_{k}\}_{k=1}^{q}\subseteq\mathbb{R}$ and $E:=\{\varepsilon^{(k)}\}_{k=1}^{q}$ with $\varepsilon^{(1)}>\cdots>\varepsilon^{(q)}$, we consider the sums
@@ 3.7.1 Hypergraph learning
We start by determining the corresponding Euler–Lagrange equations of (7). The result implies that the gradient of the energy decomposes into a sum of discrete operators, each tied to a hyperedges of a given size. The proof is in Section 4.2.1.
@@
We now study the asymptotic behavior as $n\to\infty$. The next result shows pointwise convergence of the discrete hypergraph operator to its continuum analogue.
@@
$\biggl|\left(\sum_{k=1}^{q}\lambda_{k}\Delta_{n,\varepsilon_{n}}^{(k,p)}\right)(u)(x_{i_{0}})-\left(\sum_{k=1}^{q}\lambda_{k}\Delta_{\infty}^{(k,p)}\right)(u)(x_{i_{0}})\biggr|=\mathcal{O}\left(\delta\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
@@
The proof uses a Taylor expansion of $|t|^{p-2}t$ for $p\geq 3$, relying on bounded second derivatives. For $p=2$, the expansion is exact. The details are in Section 4.2.2.
For the case $k=1$, the limiting operator $\Delta_{\infty}^{(1,p)}$ is a weighted $p$-Laplacian operator. For $k>1$, we refer to Remark 3.1 where we argue that $\Delta_{\infty}^{(k,p)}$ continues to act qualitatively like a weighted $p$-Laplacian. The main conclusion of Theorem 3.3 therefore is that hypergraph learning asymptotically behaves like a weighted variant of $p$-Laplace learning in the continuum limit. Indeed, the weighted $p$-Laplacian naturally appears as the Euler–Lagrange operator of the variational problem
(9) $\min_{v:\Omega\mapsto\mathbb{R}}\int\|\nabla v(x_{0})\|_{2}^{p}\rho(x_{0})^{2}\,\mathrm{d}x_{0},$
which defines a weighted $\mathrm{W}^{1,p}$ seminorm and serves as the continuum limit of $p$-Laplace learning on graphs, see [80].
@@
In particular, it only differs from the limiting energy of $p$-Laplacian learning (9) by a weighting of the density. This also means that, asymptotically and similarly to our conclusion from Theorem 3.3, these hypergraph models are equivalent to graph models. We thus obtain an identical characterization of well-/ill-posedness in terms of $\varepsilon_{n}$ as in [80, Theorem 2.1], now in the hypergraph context. In particular, the well-posedness, in which minimizers smoothly interpolate the known labels, is ensured if and only if $\varepsilon_{n}$ satisfies the lower bound L.2 and the upper bound $n\varepsilon_{n}^{p}\to 0$.
Similarly to [90, Remark 3.1], the above-mentioned bounds also imply that $p>d$ and we recover an intuition stemming from Sobolev spaces. Indeed, in the continuum, our functions in $\mathrm{W}^{1,p}(\Omega)$ must be at least continuous in order to satisfy pointwise constraints, i.e. be in the well-posed case: by Sobolev inequalities, this can only the case whenever $p>d$. Our results show that this condition is necessary but not sufficient as $\varepsilon_{n}$ also has to satisfy an upper bound. We also note that in practice, the condition $p>d$ often leads to $p\geq 3$, which satisfies the requirements for pointwise convergence in Theorem 3.3.
For the ill-posed case, we note that minimizers of $(\mathcal{S}\mathcal{G})_{\infty}^{(q,p)}$ are constants and therefore, for large $n$, we expect our discrete minimizers to be almost constant with spikes at the known labels (this is observed for $q=1$ in [64, 26]). The labelling problem relying on the thresholding of our minimizers is therefore rendered nonsensical, hence our denomination of ill-posed. The case $p\leq d$ is also covered by our characterization of our ill-posed case (see [90, Remark 3.3]), linking our results back to the Sobolev embedding intuition.
We now compute the Euler–Lagrange equations associated with the limiting energy functional (10), yielding the PDE:
$\sum_{k=1}^{q}\lambda_{k}\sigma_{\eta}^{(k)}\,\frac{1}{\rho(x)}\mathrm{div}\left(\rho(x)^{k+1}\left\|\nabla v(x)\right\|^{p-2}\nabla v(x)\right)=0.$
[…]
Figure 5: Pointwise and variational convergence yield distinct limiting operators, a phenomenon not previously reported for graph-based algorithms.
@@ 4 Proofs
4.1 Limiting operator $\Delta_{\infty}^{(k,p)}$
In this section, we present the proofs leading to Remark 3.1. We denote the $d$-dimensional unit sphere by $S^{d-1}$ and parametrize it by the parameters $(r,\phi_{1},\dots,\phi_{d-1})$ in spherical coordinates. We write $\mathrm{d}\Phi$ for its surface element and let $A=[0,\pi]^{d-2}\times[0,2\pi]\subset\mathbb{R}^{d-1}$ be the range of the parameters $(\phi_{1},\dots,\phi_{d-1})$. We also write $\mathrm{Area}(S^{d-1})$ for the surface area of $S^{d-1}$. Finally, in this section, $\Gamma(x)$ denotes the classical gamma function.
Lemma 4.1 (Gamma identity).
Assume that Assumption W.1 holds and let $k>1$. Then,
$C_{\mathrm{lb}}(p_{1},\dots,p_{d})\times\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma(p_{1})\cdots\Gamma(p_{d})}{\Gamma(p_{1}+\cdots+p_{d})}$
$\leq\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})|(z_{1})_{1}|^{2p_{1}-1}\cdots|(z_{1})_{d}|^{2p_{d}-1}\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}$
$\leq C_{\mathrm{ub}}(p_{1},\dots,p_{d})\times\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma(p_{1})\cdots\Gamma(p_{d})}{\Gamma(p_{1}+\cdots+p_{d})}$
where
$C_{\mathrm{ub}}(p_{1},\dots,p_{d})=\int_{[0,\infty]^{k}}\left[\prod_{s=1}^{k}\eta(r_{s})\right]\left[\prod_{j=1}^{k}\prod_{\ell=1}^{j-1}\eta(|r_{j}-r_{\ell}|)\right]r_{1}^{2(p_{1}+\cdots+p_{d}-1/2)}r_{2}^{d-1}\cdots r_{k}^{d-1}\,\mathrm{d}r_{k}\cdots\mathrm{d}r_{1}$
and
$C_{\mathrm{lb}}(p_{1},\dots,p_{d})=\int_{[0,\infty]^{k}}\left[\prod_{s=1}^{k}\eta(r_{s})\right]\left[\prod_{j=1}^{k}\prod_{\ell=1}^{j-1}\eta(|r_{j}+r_{\ell}|)\right]r_{1}^{2(p_{1}+\cdots+p_{d}-1/2)}r_{2}^{d-1}\cdots r_{k}^{d-1}\,\mathrm{d}r_{k}\cdots\mathrm{d}r_{1}.$
For $k=1$, $C_{\mathrm{ub}}(p_{1},\dots,p_{d})=C_{\mathrm{lb}}(p_{1},\dots,p_{d})$ and therefore
$\frac{2\Gamma(p_{1})\cdots\Gamma(p_{d})}{\Gamma(p_{1}+\cdots+p_{d})}\int_{0}^{\infty}\eta(r)r^{2(p_{1}+\cdots+p_{d}-1/2)}\,\mathrm{d}r=\int_{\mathbb{R}^{d}}\eta(|z|)|(z_{1})_{1}|^{2p_{1}-1}\cdots|(z_{1})_{d}|^{2p_{d}-1}\,\mathrm{d}z_{1}.$
Proof.
Let us define
$f(p_{1},\dots,p_{d};\phi_{1},\dots,\phi_{d-1})=|\cos(\phi_{1})|^{2p_{1}-1}\prod_{s=2}^{d-1}\left|\left(\prod_{\ell=1}^{s-1}\sin(\phi_{\ell})\right)\cos(\phi_{s})\right|^{2p_{s}-1}\cdot\left|\prod_{\ell=1}^{d-1}\sin(\phi_{\ell})\right|^{2p_{d}-1}.$
We recall from [14] that
$\int_{\mathbb{R}^{d}}\eta(|x|^{2})|(x)_{1}|^{2p_{1}-1}\cdots|(x)_{d}|^{2p_{d}-1}\,\mathrm{d}x=\frac{\Gamma(p_{1})\cdots\Gamma(p_{d})}{\Gamma(p_{1}+\cdots+p_{d})}\int_{0}^{\infty}\eta(t)t^{p_{1}+\cdots+p_{d}-1}\,\mathrm{d}t.$
Using spherical coordinates, with $A=[0,\pi]^{d-2}\times[0,2\pi]$, this implies that
$\int_{\mathbb{R}^{d}}\eta(|x|^{2})|(x)_{1}|^{2p_{1}-1}\cdots|(x)_{d}|^{2p_{d}-1}\,\mathrm{d}x$
$=\int_{0}^{\infty}\eta(r^{2})r^{2(p_{1}+\cdots+p_{d}-1/2)}\,\mathrm{d}r\int_{A}f(p_{1},\dots,p_{d};\phi_{1},\dots,\phi_{d-1})\,\mathrm{d}\Phi$
$=\int_{0}^{\infty}\eta(t)t^{p_{1}+\cdots+p_{d}-1}\,\mathrm{d}t\cdot\frac{1}{2}\int_{A}f(p_{1},\dots,p_{d};\phi_{1},\dots,\phi_{d-1})\,\mathrm{d}\Phi$
from which we conclude:
(11) $\frac{\Gamma(p_{1})\cdots\Gamma(p_{d})}{\Gamma(p_{1}+\cdots+p_{d})}=\frac{1}{2}\int_{A}f(p_{1},\dots,p_{d};\phi_{1},\dots,\phi_{d-1})\,\mathrm{d}\Phi.$
For each $1\leq j\leq k$, we denote by $(r_{j},\phi_{1}^{j},\dots,\phi_{d-1}^{j})$ the spherical coordinates and by $\mathrm{d}\Phi_{j}$ the corresponding surface element. We now estimate as follows:
$\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})|(z_{1})_{1}|^{2p_{1}-1}\cdots|(z_{1})_{d}|^{2p_{d}-1}\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}$
$\leq\int_{([0,\infty]\times A)^{k}}\tilde{\eta}_{\mathrm{p}}(r_{1},\dots,r_{k})r_{1}^{2(p_{1}+\cdots+p_{d}-1/2)}$
(12) $\times r_{2}^{d-1}\cdots r_{k}^{d-1}f(p_{1},\dots,p_{d};\phi_{1}^{1},\dots,\phi_{d-1}^{1})\,\mathrm{d}r_{k}\cdots\mathrm{d}r_{1}\mathrm{d}\Phi_{1}\cdots\mathrm{d}\Phi_{k}$
(13) $=C_{\mathrm{ub}}(p_{1},\dots,p_{d})\times\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma(p_{1})\cdots\Gamma(p_{d})}{\Gamma(p_{1}+\cdots+p_{d})}$
where we used the spherical change of variables $z_{j}\mapsto(r_{j},\phi_{1}^{j},\dots,\phi_{d-1}^{j})$ for $1\leq j\leq k$, the fact that $|r_{j}-r_{\ell}|\leq|z_{j}-z_{\ell}|$ and Assumption W.1 for (12) and (11) for (13).
Using the identity, $|r_{j}+r_{\ell}|\geq|z_{j}-z_{\ell}|$, we can similarly obtain
$C_{\mathrm{lb}}(p_{1},\dots,p_{d})\times\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma(p_{1})\cdots\Gamma(p_{d})}{\Gamma(p_{1}+\cdots+p_{d})}$
$\leq\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})|(z_{1})_{1}|^{2p_{1}-1}\cdots|(z_{1})_{d}|^{2p_{d}-1}\,\mathrm{d}z_{k}\cdots\mathrm{d}z_{1}.\qed$
Corollary 4.2.
Assume that Assumption W.1 holds and let $k>1$. Then,
$(p-1)\frac{\mathrm{LB}(\sigma_{\eta}^{(k,p,1)})}{\mathrm{UB}(\sigma_{\eta}^{(k,p,1)})}\leq\frac{\sigma_{\eta}^{(k,p)}}{\sigma_{\eta}^{(k,p,1)}}\leq(p-1)\frac{\mathrm{LB}(\sigma_{\eta}^{(k,p)})}{\mathrm{UB}(\sigma_{\eta}^{(k,p)})}.$
where
$\mathrm{LB}(\sigma_{\eta}^{(k,p)})=C_{\mathrm{lb}}\left(\frac{p+1}{2},\frac{1}{2},\dots,\frac{1}{2}\right)\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma\left(\frac{p+1}{2}\right)\Gamma\left(\frac{1}{2}\right)^{d-1}}{\Gamma\left(\frac{p+d}{2}\right)},$
$\mathrm{UB}(\sigma_{\eta}^{(k,p)})=C_{\mathrm{ub}}\left(\frac{p+1}{2},\frac{1}{2},\dots,\frac{1}{2}\right)\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma\left(\frac{p+1}{2}\right)\Gamma\left(\frac{1}{2}\right)^{d-1}}{\Gamma\left(\frac{p+d}{2}\right)},$
$\mathrm{LB}(\sigma_{\eta}^{(k,p,1)})=C_{\mathrm{lb}}\left(\frac{p-1}{2},\frac{1}{2},\dots,\frac{1}{2},\frac{3}{2}\right)\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma\left(\frac{p-1}{2}\right)\Gamma\left(\frac{1}{2}\right)^{d-2}\Gamma\left(\frac{3}{2}\right)}{\Gamma\left(\frac{p+d}{2}\right)}$
and
$\mathrm{UB}(\sigma_{\eta}^{(k,p,1)})=C_{\mathrm{ub}}\left(\frac{p-1}{2},\frac{1}{2},\dots,\frac{1}{2},\frac{3}{2}\right)\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma\left(\frac{p-1}{2}\right)\Gamma\left(\frac{1}{2}\right)^{d-2}\Gamma\left(\frac{3}{2}\right)}{\Gamma\left(\frac{p+d}{2}\right)}.$
For $k=1$, we have that $\mathrm{UB}(\sigma_{\eta}^{(k,p,1)})=\mathrm{LB}(\sigma_{\eta}^{(k,p,1)})$ and $\mathrm{UB}(\sigma_{\eta}^{(k,p)})=\mathrm{LB}(\sigma_{\eta}^{(k,p)})$ and therefore
$\frac{\sigma_{\eta}^{(1,p)}}{\sigma_{\eta}^{(1,p,1)}}=p-1.$
Proof.
This is just an application of Lemma 4.1. In particular,
$\mathrm{LB}(\sigma_{\eta}^{(k,p)})$ $=C_{\mathrm{lb}}\left(\frac{p+1}{2},\frac{1}{2},\dots,\frac{1}{2}\right)\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma\left(\frac{p+1}{2}\right)\Gamma\left(\frac{1}{2}\right)^{d-1}}{\Gamma\left(\frac{p+d}{2}\right)}$
$\leq\sigma_{\eta}^{(k,p)}$
$\leq C_{\mathrm{ub}}\left(\frac{p+1}{2},\frac{1}{2},\dots,\frac{1}{2}\right)\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma\left(\frac{p+1}{2}\right)\Gamma\left(\frac{1}{2}\right)^{d-1}}{\Gamma\left(\frac{p+d}{2}\right)}=\mathrm{UB}(\sigma_{\eta}^{(k,p)})$
and
$\mathrm{LB}(\sigma_{\eta}^{(k,p,1)})$ $=C_{\mathrm{lb}}\left(\frac{p-1}{2},\frac{1}{2},\dots,\frac{1}{2},\frac{3}{2}\right)\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma\left(\frac{p-1}{2}\right)\Gamma\left(\frac{1}{2}\right)^{d-2}\Gamma\left(\frac{3}{2}\right)}{\Gamma\left(\frac{p+d}{2}\right)}$
$\leq\sigma_{\eta}^{(k,p,1)}$
$\leq C_{\mathrm{ub}}\left(\frac{p-1}{2},\frac{1}{2},\dots,\frac{1}{2},\frac{3}{2}\right)\mathrm{Area}(S^{d-1})^{k-1}\frac{2\Gamma\left(\frac{p-1}{2}\right)\Gamma\left(\frac{1}{2}\right)^{d-2}\Gamma\left(\frac{3}{2}\right)}{\Gamma\left(\frac{p+d}{2}\right)}=\mathrm{UB}(\sigma_{\eta}^{(k,p,1)}).$
We note that $\Gamma(t+1)=t\Gamma(t)$ as well as
$C_{\mathrm{lb}}\left(\frac{p+1}{2},\frac{1}{2},\dots,\frac{1}{2}\right)=C_{\mathrm{lb}}\left(\frac{p-1}{2},\frac{1}{2},\dots,\frac{1}{2},\frac{3}{2}\right)$
and
$C_{\mathrm{ub}}\left(\frac{p+1}{2},\frac{1}{2},\dots,\frac{1}{2}\right)=C_{\mathrm{ub}}\left(\frac{p-1}{2},\frac{1}{2},\dots,\frac{1}{2},\frac{3}{2}\right).$
This allows us to conclude that $\mathrm{LB}(\sigma_{\eta}^{(k,p)})=(p-1)\mathrm{LB}(\sigma_{\eta}^{(k,p,1)})$ and $\mathrm{UB}(\sigma_{\eta}^{(k,p)})=(p-1)\mathrm{UB}(\sigma_{\eta}^{(k,p)})$. This leads to
$(p-1)\frac{\mathrm{LB}(\sigma_{\eta}^{(k,p,1)})}{\mathrm{UB}(\sigma_{\eta}^{(k,p,1)})}\leq\frac{\sigma_{\eta}^{(k,p)}}{\sigma_{\eta}^{(k,p,1)}}\leq(p-1)\frac{\mathrm{LB}(\sigma_{\eta}^{(k,p)})}{\mathrm{UB}(\sigma_{\eta}^{(k,p)})}.\qed$
4.2 Hypergraph learning
4.2.1 Euler-Lagrange equations

### v2

@@ 1 Introduction
    Since the dependence on $v$ in the objective function arises solely from pairwise interactions between vertices, the resulting regularization is inherently first-order. Consequently, the hypergraph itself contributes only to defining the support of interactions—that is, which and how strongly vertex pairs are connected—not to specifying the order or nature of the regularization being applied. We show in Theorems 3.2 and 3.3 that the solution of (2), in an appropriate limit, behaves as the sum of first order energies.
@@
In this paper, we study the consistency of (2) through the lens of the latter two complementary notions of convergence. Pointwise convergence allows us to show that the operator appearing in the discrete Euler–Lagrange equation of (2) converges to a weighted $p$-Laplacian operator. Variational convergence, on the other hand, enables a precise characterization of the well- and ill-posedness of hypergraph learning in the semi-supervised setting, as a function of the hypergraph construction parameters. The appropriate analytical framework for convergence of minimizers is $\Gamma$-convergence [11], which requires a common metric space in which both discrete and continuum functions can be embedded. To this end, we rely on the $\mathrm{TL}^{p}$ space introduced in [39], which is based on tools from optimal transport [88, 77]. We review these methodologies in Section 2.
@@ 1.1 Contributions
    Pointwise Convergence to Continuum Operator: We establish the quantitative pointwise convergence of the discrete Euler–Lagrange operators arising in hypergraph learning to a weighted $p$-Laplacian operator.
@@ 1.2 Related works
On the technical side, our proofs rely on nonlocal approximation results for $\mathrm{W}^{1,p}$ initially formulated in [10] and then considered under the $\Gamma$-convergence lens in [74]. The latter were extended to the discrete-to-continuum setting in [39, 84]. For other discrete-to-continuum results approximating $\mathrm{W}^{s,2}$ we refer to [26, 94, 38]. Finally, we note that our variational convergence results are purely asymptotic. Discrete-to-continuum rates have for example been obtained on similar graph learning problems in [20, 93, 28].
@@ 2.1 The $\mathrm{TL}^{p}$ Space
[39, Proposition 3.12] Let $(\mu,u)\in\mathrm{TL}^{p}$ where $\mu$ is absolutely continuous with respect to Lebesgue measure and let $\{(\mu_{n},u_{n})\}_{n=1}^{\infty}$ be a sequence in $\mathrm{TL}^{p}$. The following are equivalent:
@@
In order to compare discrete functions to continuum ones, we will let $\mu_{n}$ be the empirical measures of our samples $\{x_{i}\}_{i=1}^{n}$ and $\mu$ will be the measure from which the points are sampled. Furthermore, $\{u_{n}\}_{n=1}^{\infty}$ and $u$ will respectively be the minimizers of our discrete and continuum objectives. In order to use the above result, we need certain transport maps $T_{n}$ whose existence is guaranteed by the next theorem [43, Theorem 1.1].
@@ 3.2 Hypergraph setting
and the constant
$\sigma_{\eta}^{(k,p)}$ $=\int_{(\mathbb{R}^{d})^{k}}\tilde{\eta}_{\mathrm{p}}(\tilde{z}_{1},\dots,\tilde{z}_{k})|(\tilde{z}_{1})_{d}|^{p}\,\mathrm{d}\tilde{z}_{k}\cdots\mathrm{d}\tilde{z}_{1}.$
@@
$\Delta_{\infty}^{(k,p)}(u)(x)=\frac{\sigma_{\eta}^{(k,p)}}{2\rho(x)}\mathrm{div}(\rho(x)^{k+1}\|\nabla u\|_{2}^{p-2}\nabla u(x)).$
@@ 3.5 Variational problems for higher-order hypergraph learning
The latter set can be shown to be very closely related to the Sobolev space $\mathrm{W}^{p,2}(\Omega)$ [26, Lemma 17]. Finally, for $q\geq 1$ and positive sequences $\{\lambda_{k}\}_{k=1}^{q}\subseteq\mathbb{R}$, $P:=\{p_{k}\}_{k=1}^{q}\subseteq\mathbb{R}$ and $E:=\{\varepsilon^{(k)}\}_{k=1}^{q}$ with $\varepsilon^{(1)}>\cdots>\varepsilon^{(q)}$, we consider the sums
@@ 3.7.1 Hypergraph learning
We start by determining the corresponding Euler–Lagrange equations of (7). The result implies that the gradient of the energy decomposes into a sum of discrete operators, each tied to a hyperedges of a given size. The proof is in Section 4.1.
@@
We now study the asymptotic behavior as $n\to\infty$. The next result shows quantitative pointwise convergence of the discrete hypergraph operator to its continuum analogue.
@@
$\biggl|\left(\sum_{k=1}^{q}\lambda_{k}\Delta_{n,\varepsilon_{n}}^{(k,p)}\right)(u)(x_{i_{0}})-\rho(x_{i_{0}})\left(\sum_{k=1}^{q}\lambda_{k}\Delta_{\infty}^{(k,p)}\right)(u)(x_{i_{0}})\biggr|=\mathcal{O}\left(\delta\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
@@
The proof uses a Taylor expansion of $|t|^{p-2}t$ for $p\geq 3$, relying on bounded second derivatives. For $p=2$, the expansion is exact. The details are in Section 4.2.
@@
In particular, it only differs from the limiting energy of $p$-Laplacian learning
$\min_{v:\Omega\mapsto\mathbb{R}}\int\|\nabla v(x_{0})\|_{2}^{p}\rho(x_{0})^{2}\,\mathrm{d}x_{0}$
by a weighting of the density. This also means that, asymptotically, these hypergraph models are similar to graph models. We thus obtain an identical characterization of well-/ill-posedness in terms of $\varepsilon_{n}$ as in [84, Theorem 2.1], now in the hypergraph context. In particular, the well-posedness, in which minimizers smoothly interpolate the known labels, is ensured if and only if $\varepsilon_{n}$ satisfies the lower bound L.2 and the upper bound $n\varepsilon_{n}^{p}\to 0$.
Similarly to [94, Remark 3.1], the above-mentioned bounds also imply that $p>d$ and we recover an intuition stemming from Sobolev spaces. Indeed, in the continuum, our functions in $\mathrm{W}^{1,p}(\Omega)$ must be at least continuous in order to satisfy pointwise constraints, i.e. be in the well-posed case: by Sobolev inequalities, this can only the case whenever $p>d$. Our results show that this condition is necessary but not sufficient as $\varepsilon_{n}$ also has to satisfy an upper bound. We also note that in practice, the condition $p>d$ often leads to $p\geq 3$, which satisfies the requirements for pointwise convergence in Theorem 3.2.
For the ill-posed case, we note that minimizers of $(\mathcal{S}\mathcal{G})_{\infty}^{(q,p)}$ are constants and therefore, for large $n$, we expect our discrete minimizers to be almost constant with spikes at the known labels (this is observed for $q=1$ in [68, 27]). The labelling problem relying on the thresholding of our minimizers is therefore rendered nonsensical, hence our denomination of ill-posed. The case $p\leq d$ is also covered by our characterization of our ill-posed case (see [94, Remark 3.3]), linking our results back to the Sobolev embedding intuition.
@@ 4 Proofs
4.1 Euler-Lagrange equations of hypergraph learning